### v1

# Time to Retrain? Detecting Concept Drifts in Machine Learning Systems

Abstract
With the boom of machine learning (ML) techniques, software practitioners build ML systems to process the massive volume of streaming data for diverse software engineering tasks such as failure prediction in AIOps. Trained using historical data, such ML models encounter performance degradation caused by concept drift, i.e., data and inter-relationship (concept) changes between training and production. It is essential to use concept rift detection to monitor the deployed ML models and re-train the ML models when needed.
In this work, we explore applying state-of-the-art (SOTA) concept drift detection techniques on synthetic and real-world datasets in an industrial setting. Such an industrial setting requires minimal manual effort in labeling and maximal generality in ML model architecture. We find that current SOTA semi-supervised methods not only require significant labeling effort but also only work for certain types of ML models.
To overcome such limitations, we propose a novel model-agnostic technique (CDSeer) for detecting concept drift. Our evaluation shows that CDSeer has better precision and recall compared to the state-of-the-art while requiring significantly less manual labeling. We demonstrate the effectiveness of CDSeer at concept drift detection by evaluating it on eight datasets from different domains and use cases. Results from internal deployment of CDSeer on an industrial proprietary dataset show a 57.1% improvement in precision while using 99% fewer labels compared to the SOTA concept drift detection method. The performance is also comparable to the supervised concept drift detection method, which requires 100% of the data to be labeled. The improved performance and ease of adoption of CDSeer are valuable in making ML systems more reliable.
I Introduction
Modern systems generate huge amounts of data and are expected to reach 463 exabytes by 2025 [1] where a significant portion of the data comes in the form of time series data streams. The data can come from sensor recordings and signals from internet-of-things (IoT) and autonomous driving systems (ADS) [1]; or alerts, logs, and measurements from software systems [2, 3, 4, 5]. As these systems automatically produce data at high volume and velocity, posing a significant challenge to manual processing capabilities, practitioners develop machine learning (ML) systems powered by ML models to perform critical tasks such as predicting software failures [6].
One open challenge to the reliability of deployed ML systems is that the ML models inevitably become outdated and experience degraded performance on new data due to the constantly changing nature of real-world environments and data distributions. Such differences can be categorized as data or concept drift. Data drift occurs when production data follows a different distribution from training data or is unseen by the model during training. Conversely, concept drifts happen where the data features are the same but the relationship between the features and the class (concepts) changes. Hence, it is important to continuously test and monitor a deployed ML system for both data and concept drift. Once a drift is detected, maintainers of ML systems should be alerted that a deployed ML model is obsolete and it is time to retrain.
Current research efforts have been made to detect data drifts and concept drift detection (CDD) [7, 8, 9]. Supervised CDD is impractical as manual labeling true labels on all data is time-consuming and costly, especially when a domain expert is needed. Unsupervised CDD [10] relies on detecting data shifts and therefore cannot detect CDD efficiently. Instead, semi-supervised concept drift detection is more desired and practical due to significantly reduced manual labeling effort [8, 9]. However, such effort reduction can be affected by the frequency and the amount of data a system receives. In real-world systems, data can arrive in large volumes and at unpredictable rates, making it challenging to identify drift accurately without overwhelming computational resources or manually labeling ground truth [8, 9]. Additionally, the varying quality and representativeness of incoming data can impact the effectiveness of semi-supervised methods, necessitating strategies to adapt dynamically to changing data streams.
Particularly, we applied the SOTA semi-supervised CDD technique, i.e., ECHO [8] on various datasets and models, including one internally in Ericsson’s setting, one dataset from Google on predicting job failure [11], and four datasets from prior works. Our experiment shows that the SOTA technique requires a non-trivial amount of manual labels up to 100%, calling for better CDD techniques. When we discuss these results on real datasets and their implications if the SOTA CDDs are applied in an industrial setting, we find the limitations make the SOTA CDDs unsuitable for real-world deployment. Considering these results on real datasets and their implications for applying state-of-the-art (SOTA) concept drift detectors (CDDs) in an industrial setting, we find that the limitations make these SOTA CDDs unsuitable for real-world deployment.
We summarize three key industry concerns the SOTA techniques fail to address.
  • Excessive labeling effort. The SOTA semi-supervised CDDs require an average of 5% and in some cases, up to 100% of the time series to be labeled [12]. While this is a huge improvement compared to supervised CDDs, this is still challenging to achieve in production.
    Industry Need. Experts should be required to label as few data points as possible, and the points labeled should be as useful as possible, i.e., minimal manual labeling to achieve maximal accuracy in detecting concept drift.
  • Lack of flexibility on which data points require labeling. The SOTA CDDs request manual labels individually on demand based on metrics from datasets and models, i.e., cannot tell when and which manual labeling is required beforehand. This prevents the planning of labeling efforts in advance.
    Industry Need. Experts prefer to label multiple data points each time at their convenience. Also, experts prefer to select as few points to label as possible, preferably points selected at random so it costs them the least labeling effort.
  • Lack of generality for different ML models. The SOTA CDDs are optimized for specific model architectures and cannot be generalized for a different model architecture.
    Industry Need. There is much need for black-box CDD since the ML models can be already trained, and it is difficult to convince users to adopt a new model architecture especially if the current one is optimized for their use case.
To address the challenges above, we propose CDSeer, a novel semi-supervised CDD technique. CDSeer significantly reduces the workload for practitioners in manual labeling and is model- and distribution-agnostic. CDSeer works with ML models of different architectures and is independent of underlying data distribution. CDSeer’s innovations are as follows.
  1. 1.
    Model- and data distribution-agnostic predictable sampling for manual labeling. CDSeer utilizes a new sampling process that removes the dependence on the model architecture and data specifics. Hence CDSeer can detect concept drifts when the SOTA fails. This sampling process is also predictable in terms of the number of labels requested and the frequency of such requests, allowing for the planning of labeling efforts.
  2. 2.
    Semi-supervised learning that utilizes an inspector model to provide pseudo-GT labels for concept-drift detection. We propose a novel method to acquire pseudo labels from an inspector model to substitute true labels as the ground truth labels for concept drift prediction, which significantly lowers the number of needed labels while maintaining competitive performance to SOTA.
  3. 3.
    Model-agnostic concept drift detection. Our proposed CDSeer can be deployed to monitor ML models regardless of their architectures. Comparing this to SOTA CDD methods [9, 8] which utilizes ensemble classifiers for both prediction and concept drift detection, which is not adaptable to pre-trained industry models.
Our contributions include (1) a novel semi-supervised CDD technique that outperforms the SOTA techniques and satisfies the restrictions in industry reality; and (2) a comprehensive evaluation of CDSeer. We evaluate CDSeer on publicly available synthetic/real datasets and a proprietary dataset internal to a telecommunications company. The study on the proprietary dataset shows that CDSeer leads to comparable performance improvement on model accuracy with the supervised PHT and outperforms the SOTA semi-supervised while using very few labeled data points.
Relevance to Software Engineering in Practice. First, serving ML systems and detecting performance degradation over time (such as model obsolete) are active research areas in the Software Engineering community (e.g., [6]). Our work focuses on monitoring/testing deployed ML systems from a practical perspective, an important yet overlooked research area. Second, research on concept drift techniques can be applied to strengthen downstream SE research. For instance, ML models built to predict production system failures and performance degradation suffer from concept drifts in [6, 13]. Third, we propose a novel technique CDSeer to fill the gap between academia and software industry reality. CDSeer is shown to outperform the SOTA and is deployed in the industry. Our work is well suited to the SEIP track.
II Background
II-A Concept Drift and Data Drift
We notice the interchangeable uses of the terms “concept drift” and ”data drift” in literature can cause much confusion about their meaning [7, 10]. Hence for clarity, in this work, we define data drift and concept drift as follows.
Suppose that we have two datasets: the training $D_{training}(X,y)$ and testing $D_{testing}(X,y)$ where $X$ is the input features and $y$ is the target labels. Drifts occur when the input data distribution $P(X)$ or the distribution of the target concept of the data $P(y|X)$ changes.
  • Data drift happens when $P_{training}(X)\neq P_{testing}(X)$ and $P_{training}(y|X)$ $=$ $P_{testing}(y|X)$. Literature[14, 15, 16, 17, 18] also refer to this as “dataset shift”, “type I concept drift”, or “(virtual) concept drift” (where “virtual” is sometimes omitted). This type of drift is commonly detected via distribution monitoring or out-of-distribution detection [19, 20, 21, 22].
  • Concept drift happens when $P_{training}(X)=P_{testing}(X)$ and $P_{training}(y|X)$neq $P_{testing}(y|X)$. In literature, this is referred to as “type II concept drift”, or “real concept drift” (where “real” is sometimes omitted confusing it with data drift). This type of concept drift is the main focus of our work.
  • Concept drift and data drift may happen simultaneously when $P_{training}(X)\neq P_{testing}(X)$ and $P_{training}(y|X)$neq $P_{testing}(y|X)$. Literature names this as “type III concept drift” or just “concept drift” since it is both virtual and real at the same time. This type of drift can be detected by combining concept and data drift detectors for the best performance.
Readers can observe the mentioned variation in terminology in citations [23, 7, 24, 10]. Many methods claim concept drift detection but only monitor $P_{training}(X)\neq P_{testing}(X)$ under the assumption that concept drifts happen simultaneously with data drifts [16]. Hence, these methods cannot detect concept drift without data drifts. Our work focuses on detecting real concept drift when $P_{training}(y|X)\neq P_{testing}(y|X)$ and $P_{training}(X)=P_{testing}(X)$. We do not monitor data drift.
II-B Semi-Supervised Concept Drift Detection for Machine Learning Systems
Figure 1: High-level semi-supervised concept drift detection process.
Figure 1 illustrates our context of applying semi-supervised CDD. In our scenario, the ML models are first trained offline using labeled data. Then, they are deployed online to classify the time series. When concept drift happens, it causes accuracy deterioration, such as T3, T4, and T5 in the top right chart in Figure 1. However, in reality, the true labels of the input time series are not available. Thus, it is difficult for maintainers to observe the accuracy deterioration.
CDDs can be categorized into supervised, semi-supervised, and unsupervised. Supervised approaches require true labels for all data, which is unrealistic. Differently, unsupervised and semi-supervised approaches require true labels for zero or a few data points, thus more practical in real-world settings. The current unsupervised techniques target cases where concept and data drifts happen simultaneously, hence ineffective at detecting concept drift alone.
Based on our industry needs, we believe semi-supervised approaches are suitable to detect concept drift accurately and efficiently in production. Semi-supervised approaches require limited manual labeling and can target concept drifts alone.
In a nutshell, our work focuses on proposing a better semi-supervised approach to detect concept drifts as we believe semi-supervised approaches are better aligned with our industry needs.
III Related Works
III-A Concept Drift Detection
Supervised concept drift detection. Supervised CDD methods use the true labels of all data points to detect concept drift. They either compare statistics of the classifier error rate between the current and reference windows [25, 26, 27, 28, 29, 30] or use the intersection of confidence intervals from multiple CDDs to detect concept drift [31, 32]. While these methods are fast and accurate, they require labels for all data points which is unrealistic in production given the velocity of data streams [33].
(Semi)-supervised concept drift detection. As a more production-friendly alternative, semi-supervised CDD methods solve the problem by requiring only a subset of the data to be labeled. Existing methods can do this by using active learning [34, 35, 36], computational geometry [37], class imbalance [38], classifier confidence [39, 8], pseudo-error from a classifier ensemble [9], grid density sampling [40], ambiguity concerning the decision boundary and neighboring points [12], etc. to select only important data instances for labeling, reducing the labeling effort. However, the reliance on specific metrics to sample labels is vulnerable to data distribution and model architecture changes. Also, requiring the original model to follow certain ML architectures significantly limits adoption. We look to overcome these limitations.
Unsupervised concept drift detection. Unsupervised CDDs methods [10] detect drifts offline in batch [41, 42, 43, 44, 45, 46, 47, 48, 49, 50] or online by instances [51, 52, 53, 54, 55, 56, 57, 58, 59, 60]. These methods detect concept drift by monitoring $P_{training}(X)\neq P_{testing}(X)$ under the assumption that changes in the data distribution and the posterior probability distributions often happen simultaneously [16]. Hence, unsupervised methods cannot detect concept drift when data drift is absent (Subsection II-A).  CDSeer aims to detect real concept drift even when the input data distribution stays the same.
III-B Testing Machine Learning Systems
Testing ML systems is an important research area in SE [61], of which, black-box testing, especially of deployed models, is of particular interest due to being source code/architecture independent and generalizable to different systems.
Artificial Intelligence for IT Operations (AIOps) AIOps leverage ML models for tasks such as anomaly detection and failure prediction [62, 63, 64, 3, 65, 66, 67, 5, 2, 68] using logs, measurements, and spatial information. These models are highly susceptible to data evolution with impairs their performance and stability. Current mitigation strategies are retraining the models periodically [67, 68] or retraining the models upon alarms by supervised concept drift detectors [6]. Existing methods are very costly in manual efforts due to the amount of labeling needed. In this work, we explore CDSeer’s CDD for AIOps to reduce cost.
Model Robustness and Relevance Testing. Model testing detects mismatches between the model and data [69, 70] and ensures trained models’ resistance to data noises [71, 72, 73]. However, these techniques test pre-trained models offline and more online techniques are needed [61]. CDSeer detect concept drift online to test the robustness and relevance of ML models. Thus  CDSeer can benefit the utilization and deployment of ML models for SE tasks. In this work, we applied CDSeer to an internal dataset for anomaly detection in router systems and the Google Cluster Trace dataset for job failure prediction.
IV CDSeer: Learning-based Concept Drift Detection
In this section, we first summarize our approach at a high level (Section IV-A). Then, we describe the framework step-by-step, with design choices and implementation details.
IV-A An Overview of CDSeer
Figure 2: An overview of our proposed framework CDSeer.
Figure 2 shows an overview of our proposed concept drift detection technique namely CDSeer. In this figure, the components are separated into two parts. The grey box Online Model Prediction shows a generic data stream classification process, where the online classifier model predicts the labels of incoming streaming data points. The blue box called CDSeer summarizes our framework at a high level.
At a high level, CDSeer works as follows. As new data points arrive, CDSeer detects concept drift via an inspector model which predicts the (pseudo-) labels of the streaming data points in parallel with the online model prediction of labels. The distribution of the classification error between the online model’s predicted labels and the pseudo labels is monitored by a supervised concept drift detection module to detect the concept drifts. The window of the latest data points is then continuously clustered. Clustering allows the separation of the data points into strata where stratified samples are generated. Experts would then label these sampled points. Afterward, CDSeer stores these true labels in memory. Via semi-supervised learning, true labels in the memory are spread to the window of the newest data points, where these newly generated labels and their features are used to update the inspector model. This entire process is done online. Finally, while this is not the main focus of our experiments, we follow convention and retrain the online model when concept drift is detected using true labeled data.
IV-B Offline Training of the Inspector Model
During online deployment, the inspector model is retrained automatically (as described in later sections). However, before online deployment, the inspector model is trained offline using (a sample of) the online model’s training dataset. In this state, it cannot detect concept drifts until retrained online.
IV-C Concept Drift Detection
Data: trained $OnlineModel$, $InspectorModel$, and stream of unlabeled data points $X$
Result: the points with concept drift
1 for each $x$ in $X$ do
       /* online and insp. models predict labels */
2       $\hat{y}\leftarrow OnlineModel(x)$;
3       $\hat{y}_{insp}\leftarrow InspectorModel(x)$;
4       If $\hat{y}=\hat{y}_{insp}$ then $error\leftarrow 1$ else $error\leftarrow 0$ end;
       $CDD\leftarrow error$;
       // CDD monitors error
5       if $CDD$ raise alarm then
6             record $x$ index and notify maintainers;
7
8       end if
9      $W\leftarrow$ the latest $w$ unlabeled points including x;
10       $clusters\leftarrow$ cluster unlabeled points in $W$;
11       $candidates\leftarrow$ stratified sample from $clusters$;
12       $true\_labels\leftarrow$ labeled $candidates$ from experts;
13       if $true\_labels$ is not empty then
14             $M\leftarrow true\_labels$ and flush oldest points exceeding pre-defined fixed size;
             /* each point in $W$ is assigned a label via $LabelSpreading$ from true labels in $M$/
15             $(W,TrueLabels)\leftarrow$ spread labels from $M$ to $W$;
16             InspectorModel $\leftarrow$ trainModel$(W,TrueLabels)$;
17
18       end if
19
20 end for
Algorithm 1 Proposed Concept Drift Detection CDSeer
Each incoming unlabeled data point ($x$) in the time series is classified by both the online and the inspector models to create two labels: $\hat{y}$ and $\hat{y}_{insp}$ from respectively. The $error$ between the labels is 0 if they are the same and 1 otherwise.
$error_{i}=\begin{cases}0,&\text{if }\hat{y}=\hat{y}_{insp,i}\\ 1,&\text{otherwise}\end{cases}\forall i\in\mathcal{D}$ (1)
We utilize a CDD to monitor the distribution of the $error$s rate and alert when it crosses a threshold, signaling the online and inspector models are trained using different concepts. For this purpose, the CDD can be chosen from well-known CDDs such as the Drift Detection Method (DDM), PageHinkley Test (PHT), etc. The CDD method monitors the aggregated $error$ rate between the online and inspector models for increases surpassing a predefined threshold, which triggers a concept drift alarm between the online and the inspector models. Then, the framework records the corresponding instance and proceeds to the next data point.
IV-D Clustering and Sampling for Label Requisition
From the current data point $x$, we group the previous $w$ consecutive points including $x$ into a window $W$ which is then clustered via DBSCAN [74, 75]. DBSCAN expands from high-density core samples (measured by Euclidean distances) to form clusters, which removes the need to specify the number of clusters upfront. The cores are computed from the pre-defined minimum number of points in each cluster ($minPts$) and the maximum distance ($\epsilon$). The $minPts$ is calculated as $\frac{w}{100}$ to ensure that clusters are sufficiently large and no more than 1% of the dataset requires labeling. We estimate $\epsilon$ using a simplified version of knee detection [76]. We compute the Euclidean distance between every point in the training dataset and its nearest neighbor and plot them in ascending order. The point with the largest curvature is chosen as the $\epsilon$ since most points would have smaller distances to their neighbor and can form high-density clusters for DBSCAN.
$clusters=DBSCAN(W,Euclidean,\epsilon,\frac{w}{100})$ (2)
However, when there are no high-density cores due to randomly generated features (SEA and Sine datasets in RQ1 and RQ2), DBSCAN can form too many or only a single cluster. We address this via K-means clustering [77, 75]. The initial centroids are randomized with a count of $\frac{w}{100}$ to ensure at most 1% of the dataset requires labels. The K-means algorithm computes the Euclidean distance and assigns points closest to the centroids to their respective clusters where the mean of each cluster becomes the new centroids. This process is repeated until convergence yielding the clusters surrounding the centroids.
$clusters=KMeans(W,Euclidean,\frac{w}{100})$ (3)
When inputs are images or text (as in RQ3), we also leverage K-means for clustering after pre-processing the input. For images, the features are the pixels of the image with some caveats. First, the pixels are in grayscale and normalized to have a value between 0 and 1. Second, the size of each image needs to be very small (no more than 28x28 pixels). Datasets without these characteristics may need feature reduction and grayscale conversion. We can then compute distance treating each pixel as a feature. When the inputs are texts, we follow the original pre-processing steps to clean, trim, vectorize, and tokenize the input texts [78]. This results in the features being fixed-sized sparse feature vectors with numbered values representing the encoded vocabulary that can be used to compute distances representing similarities between texts.
Stratified random sampling data points. Via DBSCAN or K-means, the clusters from $W$ can be treated as strata to be sampled for labeling.
Expert labeling. Experts will label a subset of this sample similar to how SOTA requires experts to label the selected data points.
Rationale behind clustering. SOTA semi-supervised methods request labels based on metrics of the target data/model. For example, ECHO selects data points for labeling based on the classifier model’s confidence. Depending on the metric, SOTA may request labels for 0-100% testing data which causes the detection process to fail or be very inefficient. On the other hand, simple random label requests are inefficient and costly since we risk labeling unnecessary points. Hence, clustering is a compromise that directs the manual labeling process to the most representative points in each sample via stratified sampling.
Discussion on how CDSeer addresses industry needs. The clustering process is meant to address several industry needs. First, data points are selected so experts can label multiple points in one session. Second, the data points summarize the current data, minimizing the effort. Third, not all selected data points need to be labeled, giving the expert some choice. This allows for the flexibility needed in the industry compared to SOTA.
IV-E Propagation and Retraining the Inspector
All manually labeled data points are managed by memory $M$. $M$ is a fixed-sized queue of the newest data points with true labels available where the oldest points are automatically removed when the pre-defined size of $M$ is exceeded. Since the true labels (in $M$) are scarce, we increase their number via semi-supervised learning. To do this, we combine $M$ and $W$ and apply semi-supervised learning so true labels spread from $M$ to $W$ to get $(W,TrueLabels)$ where all points have labels. We select $LabelSpreading$ in sklearn [79], a graph inference-based algorithm that is robust to noises and can accurately spread very few labeled points to a large number of unlabeled points. This method constructs a complete weighted graph between all data points (labeled and unlabeled) where the weights are the distance between the points. Labels would spread from labeled to unlabeled data points iteratively until convergence where a higher weight slows spreading. If an unlabeled data point is assigned multiple different labels, the most frequently assigned label is selected as the final label for the point. Then, we retrain $InspectorModel$ using $(W,TrueLabels)$.
Our intuition is that when concept drift occurs, $M$ would contain a mix of both old and new concepts. When labels in $M$ are spread to $W$, it results in an inspector model that is trained on mixed concepts. We theorize that predictions from such an inspector model would increase the $error$s between the online and inspector’s predictions, triggering CDD alarms signaling the original concept has shifted. This way, the inspector’s predictions can be low-cost substitutes for the true labels. Our reasoning behind utilizing the window $W$ is two-fold: First, the true labels received are too few to train the inspector, hence, label spreading is used. The key for label spreading to work is the assumption that points belonging to the same clusters are likely to have the same labels. Second, using the window $W$ of the most recent points is likely to reflect the current concept and preserve the characteristics of the data stream. This way, supervised CDD is converted to semi-supervised, significantly reducing the number of labeled data needed.
V Evaluation Setup
This section describes the evaluation setup including computing resources, evaluation metrics, datasets, and candidate models included in this paper.
V-A Computing Environment
In RQ1 and RQ2, we experimented on a Windows computer with a Core i7-8650U processor and 32 GiB of RAM. In RQ3, to speed up the CDD experiments, the evaluations are performed on a system running Ubuntu 20.04 equipped with an AMD Ryzen 16-core processor, 256 GiB of RAM, and an NVIDIA GeForce RTX 3090 GPU.
V-B Evaluation Datasets
TABLE I: Details of the evaluation datasets. Concept drift is abbreviated as CDs.
Dataset (Source) Shape # CDs Description
RQ1 & RQ2
ROUTER (proprietary) 46,694 x 32 2 internal dataset predicting failures
Sine [26] 16,000 x 2 2 synthetic dataset
SEA [80] 16,000 x 2 2 synthetic dataset
ELEC [26] 45,312 x 9 22 real dataset predicting electricity price
NOAA [81] 18,159 x 8 3 real dataset predicting rain
GCT [11] 263,509 x 19 28 real dataset predicting job failure
RQ3
FM [82] 70,000 x 784 1 real dataset predicting clothing type
LMR [83] 50,000 x 250 1 real dataset predicting sentiment
We select eight datasets for our evaluation as shown in Table I. These publicly available datasets are collected from previous works on concept drift detection [26, 6] which consisted of three synthetic datasets (Sine, SEA, FM) and four real-world datasets (ELEC, NOAA, LMR, GCT). Also, we include one internal, real-world run-time data collected from the deployment of congestion prediction at a telecommunication company, i.e., ROUTER. In our use case, the datasets can be divided into two groups: with GT and without GT concept drift. The following datasets contain known concept drifts: Sine, SEA, FM, LMR, and ROUTER. Of these datasets, Sine, SEA, and ROUTER 100% do not have data drift. Details of the dataset are as follows:
ROUTER: An industrial and proprietary network traffic dataset containing router data from hundreds of sites over one month. This dataset includes tens of features with two features representing the delay and stability of the packet traffic having the most impact on the final class. The labels signal whether a network congestion occurs in the next hour. This dataset contains two concept drifts, where the delay and stability signal the concept drift changes due to equipment upgrades. This is an extremely imbalanced dataset where even after a 10:1 down-sampling, the positive class still consists of fewer than 4.1% of the data.
Sine [26]: A balanced, noise-free synthetic dataset with randomly generated features. For the original concept, the label is positive if the point lies below $y=sin(x)$, otherwise, it is negative. Two concept drifts are introduced in this dataset. The first is at 3,000 where the concept is the reverse of the original concept. The second is at 10,000 where the label function is changed to $y=0.5+0.3sin(3\pi x)$.
SEA [80]: The setup is similar to  Sine. For the original concept, the class is positive if the point has $att1+att2\geq 8$. Otherwise, the class is negative. The first concept drift changes the positive class to points with features $att1+att2\geq 9$ and the second concept drift changes it to points where $att1+att2\geq 7$.
Electricity (ELEC) [26]: This dataset contains data points for two months between 1996 and 1998. Each data point shows demands in different locations over a period of 30 minutes. The label identifies the change in the price related to a moving average of the last 24 hours.
NOAA [81] is a subset of the original NOAA dataset that contains 50 years of weather data from Offutt Air Force Base in Bellevue, Nebraska. The label represents whether rain was observed on each day.
Google cluster trace (GCT) [6]: first released by Google in 2011 [11]. This dataset contains the cleaned and under-sampled version of the original dataset detailing the trace data of a production cluster. The features are of the first five minutes of monitored data of the job, and the labels reflect whether the job eventually failed. This dataset is extremely imbalanced, hence, we down-sampled the non-failure class at a 3:1 ratio, leading to less than 3.20% data points that belong to the failure class.
Fashion-MNIST (FM) [82] contains images showing individual articles of clothing at low resolution (28 by 28 pixels). We reduce the number of classes to two. If the class is 0, 2, 3, 4, 6 it is a Top and the label is 0. Otherwise, it is 1. We create a synthetic concept drift at 4,000 as follows. From index 4,000, if the point’s original classes are 4 or 6 (now both are 0), it is switched to 1. In this case, it roughly translated to 366 items that have their labels changed. The concept drift detection is performed on the 10,000 images in the testing dataset.
Large Movie Review (LMR) [83] contains class balance movie reviews from the Internet Movie Database. We introduce concept drift at index 2,000 where the classes are reversed.
V-C Subject ML Models Trained Using the Datasets
We train three types of online models for the evaluation. In RQ1 and RQ2, the online models are random forest classifiers (RFC). For RQ3, two models are trained: a three-layer NN for the FM dataset and a five-layer NN for the LMR dataset.
We train the online models in RQ1 and RQ2 using the RFC implementation in scikit-learn [84]. The training data consists of the first 500/1,000 data points of the corresponding datasets including labels. The remainder of the datasets are converted into a time series without labels for testing. Hence, all experiments are conducted online where concept drift detection is active for the entire time series.
In RQ3, we evaluate the generality of CDSeer, i.e., to what extent CDSeer can detect concept drifts in different ML models as a model-agnostic approach. We trained two NNs implemented by Keras Sequential models [85, 86] using the code and data from TensorFlow [87, 78].
The first NN contains three layers for classifications of clothing images in the Fashion-MNIST dataset. The first layer flattens the input. Then, the second layer (128 nodes) learns the clothing features. Lastly, the third layer (10 nodes) outputs the classification results. This neural network is trained on the first 60,000 data points of this FM dataset.
For the textual LMR dataset, the neural network for sentiment classification consists of five layers. The first layer takes the integer-encoded text and looks up an embedding vector for each word. The third layer converts the sequence to a fixed length. The last layer is densely connected with a single output node. The second and fourth rounds are drop-out layers to reduce overfitting. This neural network is trained on the first 25,000 data points of this LMR dataset.
For all RQs, we train the inspector models using the random forest algorithm [84]. The inspector models are first trained offline using $w$ latest data points with labels in the training dataset. Then it is retrained online as described.
V-D Evaluation Metrics
We use four metrics to evaluate the performance of each concept-drift detection technique. For each CDD, we report each metric as the average of five rounds using random seed. The details of the metrics are as follows.
Model accuracy (MAcc.) (%) is the accuracy of the classification task using the online ML model with a CDD. This is computed as the percentage of correct predictions among the total data points.
Precision (%) is the percentage of true positives of all the detected concept drifts by a CDD. Recall (%) is the percentage of true positives detected by a CDD among all the GT concept drifts.
  • True positive: The number of concept drifts that are correctly detected by one CDD.
  • False positive: The number of concept drifts detected by one CDD that are incorrect and do not match with the ground-truth concept drifts.
  • False negative: The number of GT concept drifts that a CDD fails to detect.
Labels used (Lbl) (%) is the percentage of the data points that a CDD requests manual labels for.
Calculating the precision and recall when GT concept drift is unavailable. In this case, we use the result of PHT, the supervised CDD on the predictions of the original online model (no retrain) as the GT. While the concept drifts detected this way are not perfect, previous results show that they have high precision. Thus, it can reliably indicate where a subset of concept drifts are for comparison.
VI Evaluation Results
RQ1-SOTA: What are the performance and limitations of the state-of-the-art concept drift detection approaches on open-source and proprietary datasets?
Motivation. Motivated by our industry needs, we want to examine the performance of the SOTA concept-drift detection techniques on both open-source and proprietary datasets. We want to learn what their limitations are, what we can learn from them, and whether we can use the insights to devise better ideas for concept drift detection. First, we want to learn how they perform on the core functionality such as concept drift detection and the resulting accuracy. Then we want to learn how they perform on practical concerns of the concept drift detectors such as the false positive rate and the amount of labels needed.
Method. We answer this RQ using one supervised and one semi-supervised CDD approach. We set up a well-known SOTA semi-supervised concept drift detection method for comparison, i.e., ECHO [8]. It contains performance improvements over its predecessor, SAND [39], and is effective in detecting concept drift. We acquired their official Java implementation [88]. We run ECHO using their default configuration only changing the warm-up window size (to -S 1000). The default parameters are ideally their best configuration, and this change has the effect of allowing ECHO to cover up to at least 70% of the data points of each dataset. This choice comes after experimenting with various configurations, which end up with ECHO crashing when less than 30% of datasets are executed, trivializing the comparison effort. We confirmed that the results are at least on par with that observed in the original papers of ECHO and SAND [8, 39] or prior papers using them as a baseline [9, 12] on equivalent datasets.
For supervised CDD approaches, we include the Page-Hinkley test (PHT) [27] implementation in scikit-multiflow [75]. PHT is a well-known supervised CDD approach commonly used in previous work [18, 42, 57]. PHT is a variant of the Cumulative Sum [27], another well-known CDD used as the basis for concept drift detection in ECHO. As we use PHT as a supervised CDD, we allow its access to 100% of the ground-truth labels in the test datasets.
These methods are evaluated on the datasets described in Section V-B. As described in Section V-C, to evaluate PHT, we train a random forest classifier on the first 1,000 data points of the dataset with true labels to create the base model. The rest of the dataset is the test dataset. Once concept drift is detected, the base model is retrained using the most recent 1,000 data points with true labels.
TABLE II: The performance of three concept-drift detection (CDD) approaches. PHT is a supervised CDD, which requires 100% manual labeling and is unrealistic in practice. ECHO and Our Prophet are semi-supervised CDD and are more practical.
MAcc. PHT (supervised) ECHO (semi-supervised) CDSeer (semi-supervised)
w/o CDD MAcc. Prec. recall Lbl. MAcc. Prec. recall Lbl.    MAcc. Prec. recall Lbl.
ROUTER 96.56 97.88 100.0 50.0 100.0 94.60 14.3 100.0 48.1    96.76 71.4 100.0 0.6
SEA 95.91 95.93 100.0 100.0 100.0 86.44 50.0 50.0 100.0    95.55 61.5 80.0 0.6
Sine 73.69 73.63 100.0 50.0 100.0 62.25 100.0 50.0 0.03    75.97 48.1 80.0 0.6
ELEC 66.06 74.19 58.3 63.6 100.0 71.50 10.9 81.8 48.73    77.30 30.4 72.7 0.6
NOAA 76.89 77.37 66.6 66.6 100.0 77.70 6.5 100.0 100.0    77.51 15.0 80.0 0.6
Google Cloud 96.91 97.37 69.2 53.8 100.0 96.74 1.3 75.0 91.7    97.33 17.6 35.7 0.8
Results. Table II shows the results of applying PHT (a supervised CDD) and ECHO (the SOTA semi-supervised) to detect concept drifts in the six datasets. We also show the model accuracy of the original online model on the datasets for comparison. It is expected that with the help of CDD, the model accuracy will increase. Also, it is expected that the supervised CDD PHT would have a better performance than the semi-supervised CDD ECHO.
Our evaluation results show that PHT and ECHO can improve the overall model accuracy, i.e., PHT contributes to a higher model accuracy for five of the six datasets and ECHO leads to a higher or comparable model accuracy for two of the six datasets. The limited improvements of PHT and ECHO are expected for two reasons. First, the datasets contain a relatively small number of concept drifts that span a small percentage of data points. Second, these datasets contain class imbalance, hence, wrong classifications for minority classes are not as impactful. In such cases, adapting to the new concepts could cause slight decreases in accuracy. In reality, online models that take streaming data accumulated in weeks and months will have a much larger benefit from concept drift detection.
In terms of concept drift detection, our evaluation results confirm that the supervised CDD PHT achieves 100% precision in the three datasets with ground-truth concept drifts, i.e., ROUTER, SEA, and Sine. PHT fails to detect some concept drifts in ROUTER and Sine, i.e., the recall is 50%. A lower recall is expected considering the conservative design of PHT, which favors precision over recall. As a semi-supervised CDD, ECHO’s precision is lower and more unstable (ranging from 1.3% to 100%) on the six datasets compared to PHT. The recall of ECHO is comparable to that of PHT, ranging from 50% to 100%.
Another important metric of CDD is the amount of ground-truth labels required, i.e., expert labeling needs to be performed manually. A highly practical CDD is expected to have a low percentage of manual labels required. As PHT is supervised, it requires 100% of the data to be manually labeled. The semi-supervised CDD ECHO requires significantly fewer labels. However, our evaluation results find that the percentage of labels required by ECHO is unstable with a large range across different datasets, as low as 0.03% for Sine, yet an excessive amount of labeling is required for the other datasets (48.1% to 100%).
Discussions on using PHT and ECHO in industry settings. Our evaluation of PHT and ECHO exhibits significant inconsistency and raises concerns from industry partners as we cannot foresee whether the target dataset would require a lot of true labels. In our observation, the number of labels required for SOTA methods varies a lot between datasets which allows for our consistency to stand out for planning the labeling efforts. Also, as industrial partners already have pre-trained models to solve their problems, switching to another architecture with unknown results is not desirable. While PHT can work with existing models, it requires 100% data points to be labeled which is unrealistic. On the other hand, ECHO utilizes an ensemble of centroid-based k-NN models for classification and concept drift detection [8]. This type of model is not necessarily adaptable to all use cases.
RQ2-CDSeer: How does CDSeer perform compared to the SOTA methods in concept drift detection?
Motivation. As the SOTA CDD techniques do not satisfy our industry needs, we propose CDSeer including two key innovations, i.e., clustering data points and spreading true labels to reduce the excessive labeling required while improving model accuracy. In this RQ, we examine the performance of CDSeer. First, we want to evaluate the performance of CDSeer in concept drift detection using the default configuration. In addition, as CDSeer is configurable, to investigate the impacts of CDSeer’s configurations, we examine the performance of CDSeer under different configurations.
Method.
- Using CDSeer’s default configuration. We set up CDSeer to run on datasets with known concept drifts, i.e., ROUTER, SEA, and Sine. For each of the three datasets, we use the first 1,000 data points to train a random forest classifier to be the online model. In this part, we set the windows and memory sizes to our default configuration, The default window size is 1,000 and the default memory size is 15.
- Using different configurations of CDSeer. We set off to investigate how the different settings of window and memory size would impact the performance of CDSeer. In CDSeer, the window and memory size used to train the inspector model have a significant impact on the performance of concept drift detection. The window influences the clustering process and influences which data points are labeled and stored in the memory. After the labels are spread from memory, the labeled window is also the training data for the inspector model. To evaluate the impact, we vary the window and the memory size used to train the inspector model and observe the resulting performance. We experiment with memory with queue sizes 10 and 15. For the window size, we try two values of 500 and 1000, where the latter is the same size as the training data used to train the online models. Each pair of memory and window size will affect the retraining process of the inspector model, changing its prediction and the resulting concept drift detections.
As CDSeer integrates a supervised CDD to detect changes in the distribution of the prediction error between the original and the inspector models (Algorithm 1), we use PHT as the integrated supervised CDD in this experiment. Similar to RQ1, we evaluate model accuracy, precision, recall, and percentage of labels required.
Results.
- The performance of CDSeer. We list the results of CDSeer side-by-side to PHT and ECHO in Table II. Our evaluation shows that CDSeer improves the accuracy of the online models for all five datasets except for one, i.e., SEA. CDSeer has comparable performance with the supervised PHT and even a slightly higher model accuracy for three of the six datasets. Compared to the semi-supervised ECHO, CDSeer achieves higher model accuracy on all six datasets. In short, our CDSeer leads to comparable performance improvement on model accuracy with the supervised PHT and outperforms the SOTA semi-supervised ECHO.
In terms of the performance in detecting concept drifts, CDSeer outperforms the SOTA semi-supervised ECHO on both precision and recall and outperforms the supervised PHT in recall but has worse precision than PHT. For the three datasets with ground-truth labels, CDSeer has acceptable precision (more than 58.9%) and high recall (more than 90%). This gives us a fair estimation of CDSeer’s precision. For the three datasets with PHT-no-retraining labels, CDSeer has much higher precision and lower recall than ECHO. This is due to ECHO generating significantly more (false) alarms. CDSeer has lower precision than PHT, which is expected as PHT has access to 100% of the manual labels.
Concerning the amount of manual labeling needed (‘Lbl’ in Table II), the proposed CDSeer has the lowest label requirements, i.e., 0.6% to 0.8%. CDSeer requires the lowest percentage of labels compared to PHT and the SOTA semi-supervised ECHO, except for the dataset Sine. Our investigation shows that the 0.03% label requirement on Sine by ECHO is because Sine does not have many data points with low estimated confidence. Thus, this resulted in a few points meeting the criteria to be selected by ECHO. However, Sine is an exception, and for the remaining datasets, ECHO requires at least 48.1% of the data for manual labeling, which is too much to be practical. Furthermore, CDSeer requests a stable range of data for manual labels compared to the large range of data required by ECHO. The count of the labeling requests (which may request for one or more labels each) shows an even larger gap. For example, the ROUTER dataset represents one month’s worth of data, where 800 data points are logged each hour. During this period, ECHO makes 22,459 requests for data labels, while CDSeer makes 80 requests. Considering the number of labels, the number of requests, and the overhead associated with processing each request, this makes a significant difference to the experts’ labeling efforts.
Overall, CDSeer shows detection performance competitive to the state-of-the-art while utilizing fewer labeled data points. Depending on the data set, our method significantly reduces the number of labels used while maintaining comparable detection results and accuracy.
TABLE III: The performance of CDSeer under different configurations of window size and memory size.
Dataset Window Memory Precision Recall Lbl.
Size Size
ROUTER 500 10 71.4 100.0 0.9
500 15 100.0 60.0 0.7
1000 10 57.1 80.0 0.8
1000 15 71.4 100.0 0.6
Sine 500 10 58.8 100.0 0.6
500 15 50.0 70.0 0.6
1000 10 52.6 100.0 0.7
1000 15 48.1 80.0 0.6
SEA 500 10 75.0 90.0 0.7
500 15 80.0 80.0 0.7
1000 10 41.7 50.0 0.6
1000 15 61.5 80.0 0.6
- Sensitivity of CDSeer’s performance. Table III shows the performance of CDSeer with different configurations including sample size and window size. First, we look at how sample size impacts our detection results. Our results show that the best sample size varies for different datasets. For ROUTER, the sample size of 1,000 works best for the ROUTER dataset with 500 being a close second. Differently, the sample size of 500 is best for Sine and SEA. For memory size, we observe that a larger memory size may sacrifice recall for precision. For different datasets though, the best-performance memory size still varies. We conclude that memory size should be adjusted according to the size of the concept drift we want CDSeer to capture (i.e., higher recall). For concept drifts that occur only briefly, memory size should be smaller and vice versa. However, this might lead to decreased precision. Last, our evaluation shows the percentage of manual labeling required by CDSeer is stable for different window and sample size combinations. Note that with the default configuration, CDSeer still outperforms the SOTA. Our experiments confirm that CDSeer can be tuned to have a better performance.
Discussions on adopting CDSeer in industry settings. We discuss the adoption of CDSeer with our industry partners and conclude that the precision and recall of CDSeer are sufficient for industry deployment. In practice, it is preferred to have higher recall to be more alerted about potential concept drifts as they can be quickly verified to rule out false alarms. In addition, CDSeer’s extremely low manual labeling effort, especially in terms of the number of requests is desirable for planning and reducing efforts. From a practical perspective, the extra effort can be diverted to examine reported concept drift. Couple that with addressing experts’ preference of labeling multiple data points at their convenience and the flexibility in the points to be labeled makes CDSeer suitable for adoption.
CDSeer is on par with supervised CDD in terms of the accuracy of online models and outperforms the SOTA semi-supervised ECHO. CDSeer outperforms ECHO for detecting concept drifts with higher precision and recall, yet requiring significantly less manual effort for labeling. CDSeer addresses multiple industry concerns about the practicality and performance of concept drift detection, allowing deployment in industrial settings.
RQ3-Generality: Can CDSeer detect concept drift on pre-trained models?
Motivation. We followed the convention of datasets and model training in previous RQs. Yet, it remains unclear whether CDSeer can generalize to different ML models. Also, as industrial partners possess pre-trained models for their use cases, switching to other untested architecture is undesirable. While PHT can work with existing models, it requires 100% data points to be labeled which is unrealistic. On the other hand, ECHO utilizes an ensemble of centroid-based k-NN models for classification and concept drift detection [8] making it not necessarily adaptable to other use cases. Hence in this RQ, we examine CDSeer’s performance on pre-trained classifiers.
Method. We include two different Keras sequential neural networks for image classifications. The first model consists of three layers and the second model consists of five layers as described in Section V-C. We trained the two models using the released data and source code from Tensorflow document [87, 78]. Model 1 is trained using the complete 60,000 data points from the FM dataset and Model 2 is trained using the 25,000 points from the LMR dataset.
To simulate the production application, we train a random forest classifier using the last 1,000 points of the training data as the inspector model. Benefiting from more than one order of magnitude smaller training set, the inspector model is trained using significantly less time than the original model. In this RQ, we configure CDSeer to use the window size of 1,000 and the memory size of 20. Similar to prior RQs, considering the randomness of CDSeer, we run the experiments for five runs and calculate an average precision and recall.
Results. Our evaluation shows that CDSeer achieves a precision of 62.5% and a recall of 100% for Model 1 (an image classification neural network), and a precision of 80% and a recall of 80% for Model 2 (a text classification neural network). The performance of CDSeer on alternative ML models is similar to that in RQ2. As for the percentage of labels required for manual labeling, consistent with RQ2, we find CDSeer requires 0.99% and 0.98% for the two models respectively, which are around 300 to 600 data points.
CDSeer successfully detects concept drift for pre-trained models and is not limited by a model architecture (unlike ECHO) while requiring little manual labeling.
VII Threats to Validity
Threats to internal validity. CDSeer has parameters that may impact its performance. We conducted a sensitivity experiment to assess the impact of different parameters on the performance of CDSeer. Our experiment shows that CDSeer’s memory size should be proportional to the size of the concept drift for best recall. However, the default values are sufficient for most cases.
Threats to construct validity. We describe the threats below regarding propagation and the percentage of labels required. First, we use propagation to auto-label selected data points based on clustering, which may impact CDSeer’s performance. We included additional propagation methods for comparison. In particular, we observe that $LabelSpreading$ and $LabelPropagation$ models from sklearn [84] and observed similar results. Future works can investigate more propagation methods. Second, our method has inherent randomness from the sampling process. We expect to see a fluctuating number of labels required for manual labeling between runs. To address this issue, we report an average number from five runs in this paper.
Threats to external validity. Our evaluation is limited to the datasets included in this paper, and may not work for every type of concept drift. We include the commonly used and popular datasets from prior works on concept drifts. In addition, we include one proprietary dataset from industry partners and one from Google. CDSeer is model-agnostic by design. We show that CDSeer works on three models with different architectures. However, the performance of CDSeer on other models remains unknown.
VIII Conclusion
In this work, we propose CDSeer, a concept drift detection framework that can detect concept drifts with better precision and recall compared to state-of-the-art while using fewer manual labels and having fewer constraints concerning the data distribution and model architecture. CDSeer addresses industry concerns in concept drift detection and allows for its integration into existing maintenance frameworks.
Acknowledgment
This work is supported by Mitacs and Ericsson (GAIA Montreal, Canada).